# Stability theory over toroidal or Novikov-type base and canonical modifications

### v2

@@ 3.2. Higher $\Theta$-stable reduction theorem
Now we show the uniqueness assertion (2). First, note that for each positive real number $c^{\prime\prime}$, if $(\xi,c^{\prime\prime})$ is contained in a rational hypersurface, then it is defined only by $c^{\prime\prime}$ because $\xi$ is assumed to be irrational if $r>1$. We recall that from the proof of Lemma 3.11, we have $\tilde{f}(\eta_{V(\tau^{\prime})})\notin Z^{+}$ and $\lim_{t\in T,t\to p_{-\tau}}\tilde{f}(\eta_{V(\tau^{\prime}+\mathbb{R}_{\geq 0}(\vec{0},1))})\in Z^{+}$, where $p_{-\tau}$ denotes the $T$-invariant point of affine toric variety $U_{-\tau}$ over $R/\mathfrak{m}$. Here the meaning of $\lim$ is naturally justified by the extension of the $T$-action to the N-spectrum $\Delta_{-\xi}$ and leave the detailed definition to the readers. From these conditions, $c^{\prime}$ is determined canonically. Indeed, in case $c^{\prime\prime}>c^{\prime}$ and a rational polyhedral cone $\tau^{\prime\prime}$ which contains $(\xi,c^{\prime\prime})$, we would have $\tilde{f}(\eta_{V(\tau^{\prime\prime})})\in Z^{+}$ and in case $c^{\prime\prime}<c^{\prime}$ then $\lim_{t\in T,t\to p_{-\tau}}\tilde{f}(\eta_{V(\tau^{\prime\prime}+\mathbb{R}_{\geq 0}(\vec{0},1))})\notin Z^{+}$. Indeed, the latter limit is $\lim_{t\in T,t\to p_{-\tau}}\tilde{f}(\eta_{V(\tau^{\prime\prime}+\mathbb{R}_{\geq 0}(\vec{0},1))})=\tilde{f}(V(\hat{\tau}^{\prime\prime}))$, where $\hat{\tau}^{\prime\prime}$ is the cone of $\Sigma^{\prime}$ which contains $(\xi,(c^{\prime\prime},c^{\prime\prime}+\epsilon))$. However $\tilde{f}(V(\hat{\tau}^{\prime\prime}))$ is not contained in $Z^{+}$ from the higher $\Theta$-strata condition. Hence we complete the proof of the uniqueness assertion. ∎
@@
The following is a corollary to the above proof of Theorem 3.8. For that, we prepare the following, to connect to the original $\Theta$-strata ([AHLH23]).
Definition 3.12.
@@
For a higher $\Theta$-stratum of the cone $\tau$ on an algebraic stack $\mathcal{M}$, which we denote as $\mathcal{Z}^{+}\subset{\rm Map}(\Theta_{\tau},\mathcal{M})$, consider restriction map $({\rm exp}\otimes\iota)^{*}\colon\mathcal{Z}\to{\rm Map}(\Theta_{\tau},\mathcal{M})$ and suppose it gives the $\Theta$-strata structure for each $\xi$ i.e., the images are connected. 555at the moment, we are not sure how weak or strong this condition is Then we call the obtained $\Theta$-strata as restriction along $\xi$ of the $\Theta$-strata $\mathcal{Z}$ and often denote as $\mathcal{Z}_{\xi}$.
Corollary 3.13 (of proof of Theorem 3.8).
For an algebraic stack $\mathcal{M}$ with affine diagonal. We take a higher $\Theta$-stratum of the cone $\tau$ on an algebraic stack $\mathcal{M}$, which we denote as $\mathcal{Z}^{+}$ which has restrictions $\mathcal{Z}^{+}$ along each $\xi\in N\cap\tau$. As in Theorem 3.8 or [AHLH23, §6], take an arbitrary morphism $f\colon\mathop{\mathrm{Spec}}\nolimits(R)\to\mathcal{M}$ where $R$ is a valuation $\mathbbm{k}$-algebra. Suppose that $f(\mathfrak{m})$ sits in a higher $\Theta$-stratum $\mathcal{Z}^{+}$ in $\mathcal{M}$ while $f(\eta_{R})$ is not in $\mathcal{Z}^{+}$.
Then, the limit of the obtained $\Theta$-stable reduction $f|_{{\rm new},\xi}(0)$ by [AHLH23, Theorem 6.3] or Corollary 3.10 (and the one parameter family obtained $\mathcal{T}^{\prime\prime}$ in the proof of Theorem 3.8) depends only on the strata of the cone $\tau$ where $\xi$ sits in.
@@
For an algebraic stack $\mathcal{M}$ over $\mathbbm{k}$ with affine diagonal, following Notation 1, take irrational $\xi\in N_{\mathbb{R}}$, set $\Gamma_{\geq 0}:=\langle M_{\geq 0},\xi\rangle\subset\mathbb{R}$, and consider the global N-spectrum $\Delta_{\xi}^{\rm gl}=\Delta_{\xi}^{\rm gl}(\mathbbm{k})$ as in Definition 2.1. Note that the algebraic torus $T:=\mathbb{G}_{m}\otimes N$ naturally acts on $\Delta_{\xi}$.
Consider various rational polyhedral cone $\tau$ which contains $\xi$. The mapping stack from the quotient stack (a priori, fibered only over the étale site of $\mathbbm{k}$-schemes which are not necessarily finite type) $[\Delta_{\xi}^{\rm gl}/T]$ is a locally finite type algebraic stack over $\mathbbm{k}$ and has a description as follows.
@@
What we use later in the sequel papers [Od24a, Od24c] more efficiently is the following proof, rather than only the statements themselves.
@@
An analytic variant of Theorem 3.8 is as follows.
Theorem 3.15 (Analytic analogue of Theorem 3.8).
@@
Consider an arbitrary holomorphic map $\tilde{f}\colon\Delta\to M$ induces $f\colon\Delta\to\mathcal{M}:=[M/T(\mathbb{C})^{\rm an}]$ whose target is regarded as a complex analytic stack. Suppose that for a (algebraic) higher $\Theta$-stratum $\mathcal{Z}^{+}$ with ${\rm ev}_{(1,\cdots,1)}(\mathcal{Z}^{+})=Z^{+}\subset M$, $\tilde{f}^{-1}(Z^{+})\subset\Delta$ is set theoritically $\{0\}\subset\Delta$.
@@
  1. (1)
    We can modify $f$ to avoid the higher $\Theta$-stratum in the following sense. If we choose an isomorphism $\Theta^{r}\simeq\Theta_{\tau}$ for some rational regular polyhedral cone $\tau$ of a lattice $N$ of rank $r$, and fix an irrational $\xi\in\tau\setminus N_{\mathbb{Q}}$, then the map $\overline{f}$ can be taken canonically.
@@
      Each point of ${\rm Im}(\iota)$ is obtained as $\lim_{i\to\infty}t(i)\tilde{f}(s_{i})$ where $t(i)\in T(\mathbb{C})$, $s_{i}(\in\tilde{\mathbb{C}}^{*})$, $s_{i}\to 0$ for $i\to\infty$.
  2. (2)
    Such $\iota$ which satisfies (1A), (1B), (1C) is unique, up to a translation by the action of $T(\mathbb{C})$,
  3. (3)
@@
Since the proof is along the same line as that of Theorem 3.8, we only note the technical differences and the way to overcome them. The assumption that $T(\mathbb{C})\curvearrowright M$ is algebraic implies that $U_{\tau}\otimes\Delta\dashrightarrow M$ is meromorphic hence the $T$-invariant toroidal resolution of indeterminancy exists similarly as Lemma 3.11. If we apply the same arguments as Theorem 3.8 for $R:=\mathbb{C}[[t]]^{\rm conv}=\mathcal{O}^{\rm an}_{\mathbb{C},0}$ (the convergent power series ring), the rest of the arguments is almost the same except for following points:
@@
    the second item (1A) of Theorem 3.15 is slightly weaker than the algebraic version in Theorem 3.8 in that it only concerns a sequence of real numbers $\{t_{i}\}_{i}$. Thus the uniqueness statement (2) is a priori stronger.
@@
For the difference (b) can be overcome as follows. The main points are that $\mathbb{C}^{*}$ acts continuously on $M(\mathbb{C})$ with respect to their complex analytic topolgies, so that any orbit of $U(1)\otimes N$ in $M(\mathbb{C})^{\rm an}$ is compact. First we take $c^{\prime}$ and $\tau^{\prime}$ as Lemma 3.11. From the first condition of item (1B), there is some $c_{j}(j\in\{1,\cdots,l\})$ in the proof of Theorem 3.8 (before Lemma 3.11) so that the limit $\lim_{i\to\infty}\iota(t_{i})$ in the second item of (1B) is $\tilde{f}(V(\sigma_{j}))$ for the $r+1$-dimensional cone $\sigma_{j}$ of $\Sigma^{\prime}$ which intersects with $(\xi,(c_{j},c_{j}+\epsilon))$ for $0<\epsilon\ll 1$.
@@
3.3. Stable reduction along higher $\Theta$-stratification
@@
  1. (1)
    a totally ordered 666our order is converted from the notation in [HL14] for convenience to study problems related to K-stability, as in [BHLLX21] set $\Gamma$, which has a maximal element $1$,
  2. (2)
    closed777it is a priori not necessarily reduced, although in our stable reduction type theorem, the scheme structure is not really important substacks $\mathcal{M}_{<c}$ of $\mathcal{M}$ for $c\in\Gamma$ which satisfies that
    • they are monotonely increasing in the sense that we have $(\mathcal{M}_{<c})_{\rm red}\supset(\mathcal{M}_{<c^{\prime}})_{\rm red}$ if $c>c^{\prime}$ in $\Gamma$ (here, $(\mathcal{M}_{<c})_{\rm red}$ denotes the natural reduced subscheme structure on the same support),
@@
  3. (3)
    for each $c\in\Gamma$, higher $\Theta$-strata structure of type $\tau$ (Definition 3.3) for some $N,\tau$, on $\mathcal{M}_{=c}$ (we do not require to identify $N,\tau$ for different $c$)
  4. (4)
@@
Following [AHLH23, §6.2], we have the corresponding generalization as below.
Theorem 3.18.
Following the notation of Definition 3.17, we take an algebraic stack $\mathcal{M}$ of finite type with affine diagonal over $\mathbbm{k}$, and its higher $\Theta$-stratification $\{\mathcal{M}_{<c}\}_{c}$ as Definition 3.17.
Consider any $\mathbbm{k}$-morphism $f\colon\mathop{\mathrm{Spec}}\nolimits(R)\to\mathcal{M}$ for a valuation $\mathbbm{k}$-ring $(R,\mathfrak{m})$ with its fraction field $K$, which satisfies that $f(\mathop{\mathrm{Spec}}\nolimits(K))\in\mathcal{M}_{=1}$ while $f(\mathfrak{m})\in\mathcal{M}_{<1}$. Then, there is a finite sequence of modifications of $f$ and $R$ which we denote as $f_{i}\colon\mathop{\mathrm{Spec}}\nolimits(R_{i})\to\mathcal{M}(i=0,1,\cdots,l)$ with $R\to R_{i}$ such that

### v3

@@ 3.2. Higher $\Theta$-semistable reduction theorem
Now we show the uniqueness assertion (iii). First, note that for each positive real number $c^{\prime\prime}$, if $(\xi,c^{\prime\prime})$ is contained in a rational hypersurface, then it is defined only by $c^{\prime\prime}$ because $\xi$ is assumed to be irrational if $r>1$. We recall that from the proof of Lemma 3.11, we have $\tilde{f}(\eta_{V(\tau^{\prime})})\notin Z^{+}$ and $\lim_{t\in T,t\to p_{-\tau}}\tilde{f}(\eta_{V(\tau^{\prime}+\mathbb{R}_{\geq 0}(\vec{0},1))})\in Z^{+}$, where $p_{-\tau}$ denotes the $T$-invariant point of affine toric variety $U_{-\tau}$ over $R/\mathfrak{m}$. Here the meaning of $\lim$ is naturally justified by the extension of the $T$-action to the N-spectrum $\Delta_{-\xi}$ and leave the detailed definition to the readers. From these conditions, $c^{\prime}$ is determined canonically. Indeed, in case $c^{\prime\prime}>c^{\prime}$ and a rational polyhedral cone $\tau^{\prime\prime}$ which contains $(\xi,c^{\prime\prime})$, we would have $\tilde{f}(\eta_{V(\tau^{\prime\prime})})\in Z^{+}$ and in case $c^{\prime\prime}<c^{\prime}$ then $\lim_{t\in T,t\to p_{-\tau}}\tilde{f}(\eta_{V(\tau^{\prime\prime}+\mathbb{R}_{\geq 0}(\vec{0},1))})\notin Z^{+}$. Indeed, the latter limit is $\lim_{t\in T,t\to p_{-\tau}}\tilde{f}(\eta_{V(\tau^{\prime\prime}+\mathbb{R}_{\geq 0}(\vec{0},1))})=\tilde{f}(V(\hat{\tau}^{\prime\prime}))$, where $\hat{\tau}^{\prime\prime}$ is the cone of $\Sigma^{\prime}$ which contains $(\xi,(c^{\prime\prime},c^{\prime\prime}+\epsilon))$. However $\tilde{f}(V(\hat{\tau}^{\prime\prime}))$ is not contained in $Z^{+}$ from the higher $\Theta$-strata condition. Hence we complete the proof of the uniqueness assertion and hence Theorem 3.8 is now proved for the case of quotient stacks by algebraic split torus $T$.
Step 2.
In the second step, we consider the general quotient stack case i.e., when $\mathcal{M}$ is a quotient stack of the form $[M/G]$ where $M$ is an algebraic scheme over $\mathbbm{k}$ acted by an affine algebraic $\mathbbm{k}$-group $G$.
We essentially reduce the proof of the case to when $\mathcal{M}$ is a quotient stack of the form $M$ an algebraic scheme over $\mathbbm{k}$ acted by an algebraic $\mathbbm{k}$-split torus $T$ of dimension $r$.
Let us set $\mathcal{M}:=[M/G]$, where $M$ is a finite type $\mathbbm{k}$-scheme acted by an algebraic $\mathbbm{k}$-group $G$. Here we denote $m\colon T\times T\to T$ is the multiplication and $a\colon G\times M\to M$ is the action. We introduce the following concepts for the reduction to quotient stack by an algebraic split torus.
Property 1 (Local liftability).
  1. ($\mathrm{i}$)
    We call a higher $\Theta$-strata on $[Z^{+}/G]\subset\mathcal{M}$ for $G$-invariant closed subscheme $Z^{+}(\subset M)$ of type $\tau\subset N\otimes\mathbb{R}$ is (Zariski) locally liftable if the corresponding $\mathcal{Z}^{+}\subset\underline{\rm Map}(\Theta_{\tau}=[U_{\tau}/T],[M/G])$ with ${\rm ev}={\rm ev}_{(1,1,\cdots,1)}\colon\mathcal{Z}^{+}\simeq[Z^{+}/G]$ is given by the following data:
    • a disjoint union $M^{\prime}$ of Zariski open sets which covers $M$ (in particular, there is a surjective locally isomorphic morphism $c\colon M^{\prime}\twoheadrightarrow M$) with an action $a^{\prime}$ of $T$ on $M^{\prime}$ as $a^{\prime}\colon T\times M^{\prime}\to M^{\prime}$.
    • We have an algebraic $\mathbbm{k}$-morphism $M^{\prime}\to{\rm Hom}_{\rm alg.\mathbbm{k}-grp}(T,G)$ (homomorphism family), in other words $\psi\colon T\times M^{\prime}\to G$ which is homomorphism along each fiber of the second projection ${\rm pr}_{2}\colon T\times M^{\prime}\to M^{\prime}$ such that the following holds:
      1. (a)
        $c$ is $T$-equivariant with respect to $\psi$ in the sense that $a\circ c\circ{\rm pr}_{2}=c\circ a^{\prime}$ as morphisms from $T\times M^{\prime}$ to $M$.
      2. (b)
        $a^{\prime}|_{c^{-1}(Z^{+})}$ extends to $U_{\tau}\times c^{-1}(Z^{+})\to c^{-1}(Z^{+})$, which we still denote as $a^{\prime}$. Its composition with $c$ gives $\mathcal{Z}^{+}$ i.e., as a union of connected components of ${\rm Map}(\Theta_{\tau},\mathcal{M})$.
      Here, ${\rm Hom}_{\rm alg.\mathbbm{k}-grp}(T,G)$ naturally means the algebraic $\mathbbm{k}$-scheme representing homomorphisms of the algebraic $\mathbbm{k}$-groups.
  2. ($\mathrm{ii}$)
    We call a higher $\Theta$-strata on $[Z^{+}/G]\subset\mathcal{M}$ for $G$-invariant locally closed subscheme $Z^{+}(\subset M)$ of type $\tau\subset N\otimes\mathbb{R}$ is smooth (resp., fppf, étale) locally liftable if there is a smooth surjection (resp., faithfully flat surjection, étale surjection) $c\colon M^{\prime}\twoheadrightarrow M$ such that, the same holds as ($\mathrm{i}$) i.e., there exists $\psi$ which satisfies the above (a) and (b).
In practice, many examples of (higher) $\Theta$-strata satisfies the above Zariski local liftability condition ($\mathrm{i}$) and indeed we do not know any example which do not satisfies it. The main point of the above property is that it enables us the proof of Theorem 3.8 to the same arguments as Step 1 through the action $a^{\prime}$ on $M^{\prime}$. Note that this arguments do not depend on sophisticated results on algebraic stack theory, such as [AHHLR22], [AHR20, AHR19].
Now we prove the following.
Claim 4.
[…]
proof of Claim 4.
From the Białynicki-Birula decomposition [HL14, Theorem 1.4.7, 1.4.8], which extends [BB73, 4.8] (cf., e.g., also [Dri13, DG14]), the preimage of $\mathcal{Z}^{+}$ to the chart $X$, denoted as $Z^{+}(\subset X)$ satisfies that $Z^{+}/G\simeq\sqcup X^{\psi,+}/P_{\psi}$ for the attractor (blade in [HL14]) $X^{\psi,+}$ with respect to a homomorphism $\psi\colon T\to G$ in the notation of loc.cit. Note that due to the higher $\Theta$-strata condition that ${\rm ev}_{1}$ is closed immersion, for our proof, one can assume that $Z^{+}$ corresponds to only one $\psi$ as the representative of equivalence class modulo the Weyl type group $W$ of $G$. Then, from the same condition again, it follows that $Z^{+}=GX^{\psi,+}$ has a structure of closed subscheme of $X$ and there is an isomorphism $i\colon[Z^{+}/G]\simeq[X^{\psi,+}/P]$. Take the base change of $i$ by $(G\times X^{\psi,+})/P_{\psi}\to[X^{\psi,+}/P_{\psi}]$, we obtain an isomorphism  [^9] We thank T.Kinjo for the discussion on this. $(G\times X^{\psi,+})/P_{\psi}\simeq Z^{+}$. Note that $(G\times X^{\psi,+})/P_{\psi}$ (resp., $(G\times X)/P_{\psi}$ is a $X^{\psi,+}$-fiber bundle over $G/P$ (resp., $X$-fiber bundle) and in particular an algebraic space. By étale descent of the obvious fppf trivialization $(G\times X^{\psi,+})\to G\to G/P_{\psi}$ (resp., $(G\times X)\to G\to G/P_{\psi}$) of these fiber bundles, there is an étale surjection $B^{\prime}\twoheadrightarrow G/P$ so that the base change of $(G\times X)/P_{\psi}\to G/P_{\psi}$ (resp., $(G\times X^{\psi,+})/P_{\psi}\to G/P_{\psi}$ becomes trivial $X^{\psi,+}$-bundle i.e., isomorphic to product. That is,
(10) $X\times_{B^{\prime}}(G/P)$ $\simeq X\times B^{\prime},$
(11) $Z^{+}\times_{B^{\prime}}(G/P)$ $\simeq X^{\psi,+}\times B^{\prime}.$
By using $\psi\colon T\to G$, clearly there is an action on (10) which preserves its subset (11) and satisfies the desired Property 1 ($\mathrm{ii}$). ∎
Using the above Claim 4, now we can follow the arguments of Step 1. We take $M^{\prime}\twoheadrightarrow M$ as Claim 4 and consider a lift $p$ of $f(\mathfrak{m})$ for the closed point $\mathfrak{m}\in{\rm Spec}(R)$ as well as the morphism $f$. Note that since $f$ is formally étale and $R$ is complete, this is possible. Set $a^{\prime}(p_{\tau},p)=:p^{\prime}$. Then we consider the $T$-action $a^{\prime}$ on $M^{\prime}$ or a neighborhood of $p^{\prime}$ in particular, and resolve the indeterminancy of the rational map $U_{\tau}(\supset T)\times\Delta\dashrightarrow M^{\prime}$ by applying Lemma 3.11. We recall briefly what the claim of loc.cit implies: after replacing $\tau$ with a smaller enough subcone, it gives an appropriate rational polyhedral subcone of $\tau^{\prime}\subset\tau\times\mathbb{R}_{\geq 0}$ that projects isomorphically to $\tau^{\prime}$, set $\hat{\tau}^{\prime}:=\tau^{\prime}+\mathbb{R}_{\geq 0}$ and consider the corresponding affine toric variety $U_{\hat{\tau}^{\prime}}$, as well as its $1$-dimensional orbit $V(\tau^{\prime})(\simeq\mathbb{G}_{m})$ which corresponds to $\tau^{\prime}$. Then, $f$ extends to a certain $U_{\hat{\tau}^{\prime}}\times_{\mathbbm{k}[t]}\mathbbm{k}[[t]]\to M$ for an appropriate $\hat{\tau}^{\prime}$ by Lemma 3.11, and by the exactly the same reasoning (Claim 3) one can take it so that $V(\tau^{\prime})\times_{\mathbbm{k}[t]}\mathbbm{k}[[t]]$ maps outside $Z^{+}$ Note that this works by the strong condition Property 1 (b). In other words, by considering the action $a\circ(\psi\times{\rm id}_{M^{\prime}})\colon T\times M^{\prime}\to M$ and then apply Lemma 3.11 to it. Since $p^{\prime}\in M^{\prime}$, the existence of the extension of $f$ and the fact that $V(\tau^{\prime})\times_{\mathbbm{k}[t]}\mathbbm{k}[[t]]$ maps outside $Z^{+}$ for appropriate $\hat{\tau}^{\prime}$ both directly follows.
Further, taking a rational ray $\mathbb{R}_{\geq 0}(v,d)\subset\tau^{\prime}$ for a primitive integral vector $v\in\tau$ and $d\in\mathbb{Z}_{>0}$, we obtain ${\rm Spec}\mathbbm{k}[[t^{\frac{1}{d}}]]\to U_{\hat{\tau}^{\prime}}\times_{\mathbbm{k}[t]}\mathbbm{k}[[t]]$. Composing with the above $U_{\hat{\tau}^{\prime}}\times_{\mathbbm{k}[t]}\mathbbm{k}[[t]]\to M$, we obtain the required replacement $f^{\prime}$ as a desired modification.
We write the following corollary to the above proof of Theorem 3.8, which is about how the semistable reduction of [AHLH23] changes along the variation of direction $\xi$.
@@
For that, we prepare the following, to connect to the original $\Theta$-strata ([AHLH23]).
Definition 3.12 (Restriction to (usual $r=1$) $\Theta$-strata).
@@
For a higher $\Theta$-stratum of the cone $\tau$ on an algebraic stack $\mathcal{M}$, which we denote as $\mathcal{Z}^{+}\subset{\rm Map}(\Theta_{\tau},\mathcal{M})$, consider restriction map $({\rm exp}\otimes\iota)^{*}\colon\mathcal{Z}\to{\rm Map}(\Theta_{\tau},\mathcal{M})$ and suppose it gives the $\Theta$-strata structure for each $\xi$ i.e., the images are connected. Then we call the obtained $\Theta$-strata as restriction along $\xi$ of the $\Theta$-strata $\mathcal{Z}$ and often denote as $\mathcal{Z}_{\xi}$.
Corollary 3.13 (of proof of 3.8 - effect of variation of $\xi$).
Fix a quotient stack $\mathcal{M}$ over $\mathbbm{k}$. We take a higher $\Theta$-stratum of the cone $\tau$ on an algebraic stack $\mathcal{M}$, which we denote as $\mathcal{Z}^{+}$ which has restrictions $\mathcal{Z}^{+}$ along each $\xi\in N\cap\tau$. As in Theorem 3.8 or [AHLH23, §6], take an arbitrary morphism $f\colon\mathop{\mathrm{Spec}}\nolimits(R)\to\mathcal{M}$ where $R$ is a valuation $\mathbbm{k}$-algebra. Suppose that $f(\mathfrak{m})$ sits in a higher $\Theta$-stratum $\mathcal{Z}^{+}$ in $\mathcal{M}$ while $f(\eta_{R})$ is not in $\mathcal{Z}^{+}$.
Then, the limit of the obtained $\Theta$-semistable reduction $f|_{{\rm new},\xi}(0)$ by [AHLH23, Theorem 6.3] or Corollary 3.10 (and the one parameter family obtained $\mathcal{T}^{\prime\prime}$ in the proof of Theorem 3.8) varies only locally constantly along the change of $\xi$.  [^10] This phenomenon is later discussed in [BHLINK25] which appeared $8$ and a half months after our paper, for application to generalize Donaldson-Thomas type theory in a derived setup [BDINKP25]. Compare [BHLINK25, Constancy theorem, §6.1 and §7.3 case (i)].
@@
For an algebraic stack $\mathcal{M}$ over $\mathbbm{k}$ with affine diagonal, following Notation 1, take irrational $\xi\in N_{\mathbb{R}}$, set $\Gamma_{\geq 0}:=\langle M_{\geq 0},\xi\rangle\subset\mathbb{R}$, and consider the global N-spectrum $\Delta_{\xi}^{\rm gl}=\Delta_{\xi}^{\rm gl}(\mathbbm{k})$ as in Definition 2.1. Note that the algebraic split torus $T:=\mathbb{G}_{m}\otimes N$ naturally acts on $\Delta_{\xi}$.
Consider various rational polyhedral cone $\tau$ which contains $\xi$. The mapping stack from the quotient stack (a priori, fibered only over the étale site of $\mathbbm{k}$-schemes which are not necessarily finite type) $[\Delta_{\xi}^{\rm gl}/T]$ is a locally finite type algebraic stack over $\mathbbm{k}$ and has a description as follows.  [^11] See Notes added at the end of this paper, for the relation with newly appeared [BHLINK25].
@@
In the sequel papers [Od24a, Od24c], we use the idea of the following proof efficiently rather than applying the statements.
@@
An analytic version of our semistable reduction is as follows. Here we only discuss quotient stacks by algebraic tori, mainly due to the lack of enough references on analytic stacks (but cf., e.g., [Ino19a, Appendix A] and references therein) as it suffices for us for the moment.
Theorem 3.15 (Analytic analogue of Theorem 3.8 and Corollary 3.10).
@@
Consider an arbitrary holomorphic map $\tilde{f}\colon\Delta\to M$ induces $f\colon\Delta\to\mathcal{M}:=[M/T(\mathbb{C})^{\rm an}]$ whose target is regarded as a complex analytic stack. Suppose that for a (algebraic) higher $\Theta$-stratum $\mathcal{Z}^{+}$ of type $\tau$ with ${\rm ev}_{(1,\cdots,1)}(\mathcal{Z}^{+})=Z^{+}\subset M$, $\tilde{f}^{-1}(Z^{+})\subset\Delta$ is set theoritically $\{0\}\subset\Delta$.
@@
  1. ($\mathrm{i}$)
    We can modify $f$ to avoid the higher $\Theta$-stratum in the following sense. If we fix an irrational $\xi\in\tau\setminus N_{\mathbb{Q}}$, then the map $\overline{f}$ can be taken canonically.
@@
      Each point of ${\rm Im}(\iota)$ is obtained as $\lim_{i\to\infty}t(i)\tilde{f}(s_{i})$ where $t(i)\in T(\mathbb{C})$, $s_{i}(\in\Delta)$, $s_{i}\to 0$ for $i\to\infty$.
  2. ($\mathrm{ii}$)
    Such $\iota$ which satisfies ($\mathrm{i}$B), ($\mathrm{i}$C) is unique, up to a translation by the action of $T(\mathbb{C})$,
  3. ($\mathrm{iii}$)
@@
Since the proof is along the same line as that of Step 1 of Theorem 3.8, we only note the technical differences and the way to overcome them. The assumption that $T(\mathbb{C})\curvearrowright M$ is algebraic implies that $U_{\tau}\otimes\Delta\dashrightarrow M$ is meromorphic hence the $T$-invariant toroidal resolution of indeterminancy exists similarly as Lemma 3.11. If we apply the same arguments as Theorem 3.8 for $R:=\mathbb{C}[[t]]^{\rm conv}=\mathcal{O}^{\rm an}_{\mathbb{C},0}$ (the convergent power series ring), the rest of the arguments is almost the same except for following points:
@@
    the item ($\mathrm{i}$C) of Theorem 3.15 is slightly weaker than the algebraic version in Theorem 3.8 in that it only concerns a sequence of real numbers $\{t_{i}\}_{i}$. Thus the uniqueness statement ($\mathrm{ii}$) is a priori stronger.
@@
For the difference (b) can be overcome as follows. The main points are that $\mathbb{C}^{*}$ acts continuously on $M(\mathbb{C})$ with respect to their complex analytic topolgies, so that any orbit of $U(1)\otimes N$ in $M(\mathbb{C})^{\rm an}$ is compact. First we take $c^{\prime}$ and $\tau^{\prime}$ as Lemma 3.11. From the first condition of item ($\mathrm{i}$B), there is some $c_{j}(j\in\{1,\cdots,l\})$ in the proof of Theorem 3.8 (before Lemma 3.11) so that the limit $\lim_{i\to\infty}\iota(t_{i})$ in the second item of ($\mathrm{i}$B) is $\tilde{f}(V(\sigma_{j}))$ for the $r+1$-dimensional cone $\sigma_{j}$ of $\Sigma^{\prime}$ which intersects with $(\xi,(c_{j},c_{j}+\epsilon))$ for $0<\epsilon\ll 1$.
@@
3.3. Semistable reduction along higher $\Theta$-stratification: by repetition of Corollary 3.10
@@
  1. ($\mathrm{i}$)
    a totally ordered 121212our order is converted from the notation in [HL14] for convenience to study problems related to K-stability, as in [BHLLX21] set $\Gamma$, which has a maximal element $1$,
  2. ($\mathrm{ii}$)
    closed131313it is a priori not necessarily reduced, although in our semistable reduction type theorem, the scheme structure is not really important substacks $\mathcal{M}_{\leq c}$ of $\mathcal{M}$ for $c\in\Gamma$ which satisfies that
    • they are monotonely increasing in the sense that we have $(\mathcal{M}_{\leq c})_{\rm red}\supset(\mathcal{M}_{\leq c^{\prime}})_{\rm red}$ if $c>c^{\prime}$ in $\Gamma$ (here, $(\mathcal{M}_{\leq c})_{\rm red}$ denotes the natural reduced subscheme structure on the same support),
@@
  3. ($\mathrm{iii}$)
    We write the open substack $\mathcal{M}_{\leq c}\setminus{\rm ev}_{1}(\mathcal{M}_{=c})$ as $\mathcal{M}_{<c}$, which we assume to coinside with $\bigcup_{c^{\prime}<c}\mathcal{M}_{\leq c^{\prime}}$, and we sometimes write ${\rm ev}_{1}(\mathcal{M}_{=c})\subset\mathcal{M}$ simply as $\mathcal{M}_{=c}$ if there is no possibility of confusion.
    Then, for each $c\in\Gamma$, there is a higher $\Theta$-strata structure of type $\tau$ (Definition 3.3) for some $N,\tau$, which we write as ${\rm ev}_{1}\colon\mathcal{M}_{=c}\hookrightarrow\mathcal{M}_{\geq c}$. Here, we do not require to identify $N,\tau$ for different $c$ and we allow any rational polyhedral $\tau$ for each $N,c$.
  4. ($\mathrm{iv}$)
@@
By repeating Theorem 3.8 and Corollary 3.10 for each higher $\Theta$-strata, we obtain their following version.
Theorem 3.18 (Semistable reduction along higher $\Theta$-stratification).
Following the notation of Definition 3.17, we take an algebraic stack $\mathcal{M}$ which is a quotient stack over $\mathbbm{k}$, and its higher $\Theta$-stratification $\{\mathcal{M}_{<c}\}_{c}$ as Definition 3.17.
Consider any $\mathbbm{k}$-morphism $f\colon\mathop{\mathrm{Spec}}\nolimits(R)\to\mathcal{M}$ for a valuation $\mathbbm{k}$-ring $(R,\mathfrak{m})$ with its fraction field $K$, which satisfies that $f(\mathop{\mathrm{Spec}}\nolimits(K))\in\mathcal{M}_{=1}$ while $f(\mathfrak{m})\in\mathcal{M}_{<1}$. Then, there is a finite sequence of modifications of $f$ and $R$ which we denote as
$f_{i}\colon\mathop{\mathrm{Spec}}\nolimits(R_{i})\to\mathcal{M}(i=0,1,\cdots,l)$
with $R\to R_{i}$ such that